# LMU-ASC 06/24

### v1

@@ Abstract
We implement Genetic Algorithms for triangulations of four-dimensional reflexive polytopes which induce Calabi-Yau threefold hypersurfaces via Batryev’s construction. We demonstrate that such algorithms efficiently optimize physical observables such as axion decay constants or axion-photon couplings in string theory compactifications. For our implementation, we choose a parameterization of triangulations that yields homotopy inequivalent Calabi-Yau threefolds by extending fine, regular triangulations of two-faces, thereby eliminating exponentially large redundancy factors in the map from polytope triangulations to Calabi-Yau hypersurfaces. In particular, we discuss how this encoding renders the entire Kreuzer-Skarke list amenable to a variety of optimization strategies, including but not limited to Genetic Algorithms. To achieve optimal performance, we tune the hyperparameters of our Genetic Algorithm using Bayesian optimization. We find that our implementation vastly outperforms other sampling and optimization strategies like Markov Chain Monte Carlo or Simulated Annealing. Finally, we showcase that our Genetic Algorithm efficiently performs optimization even for the maximal polytope with Hodge numbers $h^{1,1}=491$, where we use it to maximize axion-photon couplings.
May 14, 2024
@@ 2.1 Generating two-face inequivalent triangulations
For such an encoding to be most useful, one needs an efficient method to map DNA $\mathcal{C}$ to the associated FRSTs $\mathcal{T}$ of $\Delta^{\circ}$. In practice, the algorithm developed by one of the authors [43] achieves precisely that. Specifically, instead of constructing the full set of FRSTs first and then modding out by two-face inequivalence, [^3] For $h^{1,1}\geq 10$, this is usually computationally infeasible and even at lower values of $h^{1,1}$ requires huge memory resources. For instance, computing all FRSTs for the unique polytope with $(h^{1,1},h^{1,2})=(9,19)$ required more than several hundreds of GBs of memory leading to #$\#\text{FRSTs}=162917$, while directly generating the only 14 NTFE FRSTs using the methods described here required $4$MB. one can efficiently construct a polyhedral cone for which any vector in its strict interior generates the desired two-face restrictions $\mathcal{C}$. To be more precise, this cone is the intersection of the secondary cone of each two-face triangulation, embedded in the height space of the entire polytope. Any vector in the strict interior of this intersection is also in the strict interior of each two-face cone. Thus, it defines a triangulation $\mathcal{T}$ with the imposed two-face restrictions, $\mathcal{C}$. In this way, two-face inequivalent triangulations are generated much more time and memory efficiently by requiring orders of magnitude fewer operations.
For our purposes, we need to understand which choices of two-face triangulations $\mathcal{C}$ are realizable as a regular triangulation $\mathcal{T}$ of the ambient polytope. This is equivalent to asking whether the intersection of the secondary cones is full-dimensional. This is achieved by constructing the constraint programming problem of finding any point, $x$, satisfying $Hx>0$, where $H$ are the inwards-facing hyperplane normals of the intersected cone. If no such $x$ exists, then no triangulation of the polytope has said two-face restrictions $\mathcal{C}$ [^4] We recall that only height vectors in the strict interior of the secondary cone give rise to the desired triangulation.
@@ 2.2 Equivalences beyond two-face triangulations
#$\#\text{FRSTs}\lesssim 1.53\times 10^{928}\,,\quad\#\text{NTFE FRSTs}\lesssim 1.65\times 10^{428}\,.$ (2.2)
@@ 3.2 Implementation for polytope triangulations
    • rank selection (RS): we assign ranks $r$ to each individual based on their fitness from which define again a linear probability distribution $1-r/P$ to select individuals for. In this way, a more balanced selection process can be ensured.
@@
    Population generation: After constructing the new two-face triangulations, we check whether they extend to a full FRST of the polytope $\Delta^{\circ}$ as outlined above (recall §2.1). We discard those DNA which fail to extend to an FRST, while the remaining ones are used to construct a new population. We additionally enforce that DNA are pairwise distinct in the new population.
@@ 3.3 Hyperparameter tuning from Bayesian optimization
Beyond modeling $f(x)$ as the means of the posterior distribution of $g(x)$, this statistical approach sheds light on the optimal points to sample next: in general, we want to balance exploitation (sampling near previously observed) and exploration (sampling points far from previously sampled points), and the means of our process reveal where the function is largest (where we ought to exploit) while the standard deviations establish where we understand $f$ the least (where we ought to explore). In practice, a user intructs a Bayesian Optimization algorithm to sample $N$ points in the domain: the first $M$ are sampled randomly and then the subsequent $N-M$ are chosen by maximizing [^15] Let us stress that acquisition functions are nicer and quicker to evaluate than the expensive $f$, and are thus much more amenable to simpler optimization methods. so-called acquisition functions, such as the expected improvement of the largest functional value sampled thus far. For further discussion on acquisition functions, see e.g. figures $1.2-1.4$ and the surrounding text in [66]. Upon evaluating $f$ at these $x_{1},\dots,x_{N}$, the algorithm then returns $\mathrm{argmax}_{x_{i}}f$, the $x_{i}$ corresponding to the largest seen value of $f$.
@@ 4.1 Maximizing the volume — $h^{1,1}=23$
This amounts to $3317764^{4}\times 6^{4}=$ 331776 sets of two-face FRTs, i.e., choices of DNA. By direct computation, we verified that the number of NTFE FRSTs is
@@
Let us start with a simple task to showcase the performance of our GA. We pick the volume $\mathcal{V}$ of the Calabi-Yau threefold $X^{\circ}$ as our optimization target: the task will be to maximize it. To compute it, we choose a basis $\{\omega_{i}\}_{i=1}^{h^{1,1}(X^{\circ})}$ of $H^{2}(X^{\circ},\mathbb{Z})$ and define the Kähler class $J=\sum_{i}t^{i}\,\omega_{i}$ of $X^{\circ}$ in terms of Kähler parameters $\{t^{i}\}_{i=1}^{h^{1,1}(X^{\circ})}$. The volume of $X^{\circ}$ is then given by
@@
In particular, we identify that the two largest values of $\log_{10}(\mathcal{V})$ ($6.737$ and $6.713$) have associated DNA separated by a Hamming distance of $5$ (with the maximal Hamming distance being $8$ because the polytope has $8$ non-trivial two-faces). If we normalize the feature distance and Hamming distances by the ranges of their respective distributions, the feature distance is 100$\sim$ 100$\times$ 100 × smaller, so we decide to treat these two DNA as degenerate global maxima and probe the number of unique encountered DNA $D$ required to find either one of them. In Fig. 5 we present a violin plot comparing the distribution of $D$ for our two best models from the previous plot: the genetic algorithm and best-first search. [^23] We stress that the GA used for Fig. 4 were designed to quickly locate the tail of the $\log_{10}(\mathcal{V})$ distribution but that their low genetic diversity (having a population size of $10$) and low mutation rate obstructs them from exploring the entire search space, as is necessary to find global extrema. To combat this, to produce Fig. 5 we employ an ad hoc prescription where we increase the population size and mutation rate at two points in the training process to facilitate exploration. In particular, after 1000 unique evaluated DNA we increase the mutation rate and population size to $0.1$ and $100$, respectively, then after $4000$ unique evaluated DNA we increase to $0.2$ and $200$, respectively. We see that the two methods perform comparably, with the GA having heavier tails on both ends than BeFS. Importantly, for both models one of the two best DNA are encountered with $\leq 16000$ unique DNA evaluated, meaning global extrema are reliably found with exposure to less than $5\%$ of the full search space.
To summarize, we established for the first time that GAs can tackle optimization problems defined on the space of NTFE FRSTs or, equivalently, on the space of homotopy inequivalent Calabi-Yau hypersurfaces. What is more, we provided evidence that the notion of DNA for such manifolds as introduced in §2.1 is a useful encoding not just for GAs, but more generally also for other optimization algorithms.242424 For example, the utility of the two-face parameterization of triangulations in the context of Reinforcement Learning will be featured in the upcoming article [71].
@@ 4.2 Optimizing axion decay constants — $h^{1,1}=60$
In model building in string cosmology, one would like the value of $f$ to take values in a particular range. For instance, fuzzy dark matter usually requires $f\approx 10^{16}$ GeV (see e.g [80]) plus additional constraints on the mass of the axion. In the optimization language, this corresponds to solving the inverse problem: for a given target value $f_{*}$, which NTFE FRSTs $\mathcal{T}$ give rise to decay constants $f(\mathcal{T},t_{\star}^{a})\approx f_{*}$? Below, we demonstrate that such problems can be addressed with our GA. For our search, we use $f_{*}=10^{14}$eV as the target.
@@
#NTFE $\#\text{NTFE FRSTs}\leq 4^{6}\times 64^{5}\times 168^{3}\times 734^{6}\approx 3.3\times 10^{36}\,.$ (4.10)
@@
Lastly, let us highlight the GA’s effectiveness in finding novel NTFE FRSTs compared to random sampling of two-face FRTs. In Fig. 8, we plot the number of “failed” extensions of configurations of two-face FRTs as a function of the generation. Let us reiterate that extensions “fail” whenever the FRTs of two-faces cannot be consistently extended to a regular triangulation of $\Delta^{\circ}$. At generation 00, we randomly sample FRTs of two-faces and need $4999$ choices $\mathcal{C}$ of two-face triangulations to obtain the desired $P=100$ NTFE FRSTs. As becomes evident from Fig. 8, the number of such unextandable choices of two-face triangulations decreases rapidly across the entire duration of the search. Said differently, the GA develops a strategy to find new configurations of two-face triangulations that give rise to novel FRSTs. In fact, over the entire space of states visited by the GA, only $8.4\%$ of FRSTs are redundant, i.e., appear more than once.
@@ 4.3 Maximizing axion-photon couplings — $h^{1,1}=491$
As a final application of our GA, we study the largest polytope in the KS database with $(h^{1,1},h^{1,2})=(491,11)$, whose associated Calabi-Yau geometries are expected to dominate the entire KS set. Indeed, according to [37], the search space of NTFE FRSTs of this polytope is bounded by [^26] The original bound of #$\#\text{CYs}\lesssim 1.65\times 10^{428}$ obtained in [37] by estimating the number of two-face FRTs can be further improved to #$\#\text{CYs}<10^{296.1}$ [82] by exactly counting and using more stringent bounds on the number of two-face FRTs of the polytope with $h^{1,1}=491$. #$\#\text{CYs}\lesssim 1.65\times 10^{428}$. Because this polytope has both the largest search space in KS and features the most complex and computationally intensive geometries in this database, it represents an important and exciting target for optimization. Thanks to recent advances in computational geometry [6] rendering quantities such as triple intersection numbers tractable even at $h^{1,1}=491$, targeting this polytope with optimization techniques is now within reach.
@@
For the two-faces with $17,13,7$ and $3$ points, we find respectively $19594$, $204$, $5$ and $1$ FRTs. There are three two-faces with more than $100$ points for which it is currently infeasible to compute the full set of FRTs. We explored two approaches to these two-faces: the simplest was merely to choose a Delaunay triangulation for each one, and allow the GA to explore the two-face triangulations for all other two-faces. We additionally investigated the use of random subsamples of the triangulations for these two-faces (using methods such as the ‘fast’ algorithm of [37]), but found that this decreased extendability substantially, so we decided to use the former approach and defer a more comprehensive study of the DNA at $h^{1,1}=491$ for future work. In summary, then, the GA’s search space had size
#NTFE FRSTs accessible to $\#\text{NTFE FRSTs accessible to GA}\leq 5^{2}\times 204^{2}\times 19594^{2}\approx 3.99\times 10^{14}\,.$ (4.13)
@@
Figure 9: Best $\|g_{a\gamma\gamma}\|$ found as a function of number of unique DNA encountered while optimizing, averaged over $17$ GA runs and compared with average performance of random DNA sampling. The average maximum value of $\log_{10}(\|g_{a\gamma\gamma}\|/\text{GeV}^{-1})$ after 1500 unique evaluated DNA is $-9.73\pm 0.25$ for the GA, and $-10.74\pm 0.11$ for random sampling for $95\%$ confidence.
@@
Our optimization task is to maximize an effective overall axion photon coupling, given by the Euclidean norm of the vector of axion-photon couplings for each of the $h^{1,1}=491$ axions in our theory. We denote this $\|g_{a\gamma\gamma}\|$. For this to be well-defined, we must choose a divisor to host QED and a point in the Kähler cone: to pick these, we take some moderate inspiration from phenomenology. In particular, if we assume QCD and QED are hosted on separate divisors, the correct IR couplings are realized for divisor volumes of $40$ and $127.5$, respectively. We decide to compute the tip of the stretched Kahler cone and a basis of the smallest prime toric divisors, dilate until the smallest basis prime toric divisor $D$ has volume $40$ (thinking of this as our QCD divisor), and then place QED on the next smallest basis divisor $D^{\prime}$ which intersects $D$, requiring that this divisor have volume less than $127.5$. This determines our point in moduli space and the QED divisor: if such a dilation cannot be performed, or such a QED candidate divisor does not exist, we set the target to be $\infty$- ∞. In this way, our $g_{a\gamma\gamma}$ computation is mildly phenomenologically informed, but we stress that our optimized values should taken with a grain of salt: we omit many crucial checks of the legitimacy of the associated low-energy EFT.
@@
We train GAs with target $\log_{10}\|g_{a\gamma\gamma}\|$ with population size $P=50$ until each has evaluated $1500$ unique DNA. It takes $\mathcal{O}(\text{hours})$ on standard hardware for a single run, rendering our hyperparameter optimization methods — for which we were performing hundreds of runs — impractical. We compare the outcome of the GA with random sampling with error bars from averaging over $17$ runs in Fig. 9. The GA distinguishes itself from a brute force search approach after exposure to only $\sim 200$ unique DNA from the search space and continues to improve at a much faster rate than random sampling. In particular, within $1500$ unique seen DNA the GA has found effective axion-photon couplings over an order of magnitude larger than encountered by random sampling. To make this more quantitative, the average maximum value of $\|g_{a\gamma\gamma}\|$ found by the two methods after 1500 unique evaluated DNA is
@@
$\log_{10}(\|g_{a\gamma\gamma}\|/\text{GeV}^{-1})_{\text{max}}=\begin{cases}-9.73\pm 0.25&\text{Genetic Algorithm}\,,\\ -10.74\pm 0.11&\text{Random Sampling}\end{cases}$ (4.16)
@@
for $95\%$ confidence. Thus, we find the encouraging result that, even without any kind of tuning or hyperparameter optimization, a GA can substantially improve upon brute-force searching even for the polytope with the largest and most computationally intricate search space.

### v2

@@ Abstract
We implement Genetic Algorithms for triangulations of four-dimensional reflexive polytopes which induce Calabi-Yau threefold hypersurfaces via Batyrev’s construction. We demonstrate that such algorithms efficiently optimize physical observables such as axion decay constants or axion-photon couplings in string theory compactifications. For our implementation, we choose a parameterization of triangulations that yields homotopy inequivalent Calabi-Yau threefolds by extending fine, regular triangulations of two-faces, thereby eliminating exponentially large redundancy factors in the map from polytope triangulations to Calabi-Yau hypersurfaces. In particular, we discuss how this encoding renders the entire Kreuzer-Skarke list amenable to a variety of optimization strategies, including but not limited to Genetic Algorithms. To achieve optimal performance, we tune the hyperparameters of our Genetic Algorithm using Bayesian optimization. We find that our implementation vastly outperforms other sampling and optimization strategies like Markov Chain Monte Carlo or Simulated Annealing. Finally, we showcase that our Genetic Algorithm efficiently performs optimization even for the maximal polytope with Hodge numbers $h^{1,1}=491$, where we use it to maximize axion-photon couplings.
October 27, 2025
@@ 2.1 Generating two-face inequivalent triangulations
For such an encoding to be most useful, one needs an efficient method to map DNA $\mathcal{C}$ to the associated FRSTs $\mathcal{T}$ of $\Delta^{\circ}$. In practice, the algorithm developed by one of the authors [43] achieves precisely that. Specifically, instead of constructing the full set of FRSTs first and then modding out by two-face inequivalence, [^3] For $h^{1,1}\geq 10$, this is usually computationally infeasible and even at lower values of $h^{1,1}$ requires huge memory resources. For instance, computing all FRSTs for the unique polytope with $(h^{1,1},h^{1,2})=(9,19)$ required more than several hundreds of GBs of memory leading to $\text{FRSTs}=162917$, while directly generating the only 14 NTFE FRSTs using the methods described here required 44MB. one can efficiently construct a polyhedral cone for which any vector in its strict interior generates the desired two-face restrictions $\mathcal{C}$. To be more precise, this cone is the intersection of the secondary cone of each two-face triangulation, embedded in the height space of the entire polytope. Any vector in the strict interior of this intersection is also in the strict interior of each two-face cone. Thus, it defines a triangulation $\mathcal{T}$ with the imposed two-face restrictions, $\mathcal{C}$. In this way, two-face inequivalent triangulations are generated much more time and memory efficiently by requiring orders of magnitude fewer operations.
For our purposes, we need to understand which choices of two-face triangulations $\mathcal{C}$ are realizable as a regular triangulation $\mathcal{T}$ of the ambient polytope. This is equivalent to asking whether the intersection of the secondary cones is full-dimensional. This is achieved by constructing the constraint programming problem of finding any point, $x$, satisfying $Hx>0$, where $H$ are the inwards-facing hyperplane normals of the intersected cone. If no such $x$ exists, then no triangulation of the polytope has said two-face restrictions $\mathcal{C}$. [^4] We recall that only height vectors in the strict interior of the secondary cone give rise to the desired triangulation.
@@ 2.2 Equivalences beyond two-face triangulations
$\text{FRSTs}\lesssim 1.53\times 10^{928}\,,\quad\#\text{NTFE FRSTs}\lesssim 1.65\times 10^{428}\,.$ (2.2)
@@ 3.2 Implementation for polytope triangulations
    • rank selection (RS): we assign ranks $r$ to each individual based on their fitness from which define again a linear probability distribution 1r/P1-r/P to select individuals for. In this way, a more balanced selection process can be ensured.
@@
    Population generation: After constructing the new two-face triangulations, we check whether they extend to a full FRST of the polytope $\Delta^{\circ}$ as outlined above (recall §2.1). We discard those DNA which fail to extend to an FRST, while the remaining ones are used to construct a new population. In particular, we achieve this by assigning a target of -$\infty$ to such DNA, which precludes their selection in the GA. We additionally enforce that DNA are pairwise distinct in the new population.
@@ 3.3 Hyperparameter tuning from Bayesian optimization
Beyond modeling $f(x)$ as the means of the posterior distribution of $g(x)$, this statistical approach sheds light on the optimal points to sample next: in general, we want to balance exploitation (sampling near previously observed) and exploration (sampling points far from previously sampled points), and the means of our process reveal where the function is largest (where we ought to exploit) while the standard deviations establish where we understand $f$ the least (where we ought to explore). In practice, a user intructs a Bayesian Optimization algorithm to sample $N$ points in the domain: the first $M$ are sampled randomly and then the subsequent NMN-M are chosen by maximizing [^15] Let us stress that acquisition functions are nicer and quicker to evaluate than the expensive $f$, and are thus much more amenable to simpler optimization methods. so-called acquisition functions, such as the expected improvement of the largest functional value sampled thus far. For further discussion on acquisition functions, see e.g. figures 1.21.41.2-1.4 and the surrounding text in [66]. Upon evaluating $f$ at these $x_{1},\dots,x_{N}$, the algorithm then returns $\mathrm{argmax}_{x_{i}}f$, the $x_{i}$ corresponding to the largest seen value of $f$.
@@ 4.1 Maximizing the volume — $h^{1,1}=23$
This amounts to $4^{4}\times 6^{4}=331776$ sets of two-face FRTs, i.e., choices of DNA. By direct computation, we verified that the number of NTFE FRSTs is
$\text{NTFE FRSTs}=331192\leq 331776\,.$ (4.3)
@@
In particular, we identify that the two largest values of $\log_{10}(\mathcal{V})$ ($6.737$ and $6.713$) have associated DNA separated by a Hamming distance of $5$ (with the maximal Hamming distance being $8$ because the polytope has $8$ non-trivial two-faces). If we normalize the feature distance and Hamming distances by the ranges of their respective distributions, the feature distance is $\sim 100\times$ smaller, so we decide to treat these two DNA as degenerate global maxima and probe the number of unique encountered DNA $D$ required to find either one of them. In Fig. 5 we present a violin plot comparing the distribution of $D$ for our two best models from the previous plot: the genetic algorithm and best-first search. [^23] We stress that the GA used for Fig. 4 were designed to quickly locate the tail of the $\log_{10}(\mathcal{V})$ distribution but that their low genetic diversity (having a population size of $10$) and low mutation rate obstructs them from exploring the entire search space, as is necessary to find global extrema. To combat this, to produce Fig. 5 we employ an ad hoc prescription where we increase the population size and mutation rate at two points in the training process to facilitate exploration. In particular, after 1000 unique evaluated DNA we increase the mutation rate and population size to $0.1$ and $100$, respectively, then after $4000$ unique evaluated DNA we increase to $0.2$ and $200$, respectively. We see that the two methods perform comparably, with the GA having heavier tails on both ends than BeFS. Importantly, for both models one of the two best DNA are encountered with $\leq 16000$ unique DNA evaluated, meaning global extrema are reliably found with exposure to less than $5\%$ of the full search space.
To summarize, we established for the first time that GAs can tackle optimization problems defined on the space of NTFE FRSTs or, equivalently, on the space of homotopy inequivalent Calabi-Yau hypersurfaces. What is more, we provided evidence that the notion of DNA for such manifolds as introduced in §2.1 is a useful encoding not just for GAs, but more generally also for other optimization algorithms. [^24] For example, the utility of the two-face parameterization of triangulations in the context of Reinforcement Learning will be featured in the upcoming article [71].
@@ 4.2 Optimizing axion decay constants — $h^{1,1}=60$
In model building in string cosmology, one would like the value of $f$ to take values in a particular range. For instance, fuzzy dark matter usually requires $f\approx 10^{16}$ GeV (see e.g [80]) plus additional constraints on the mass of the axion. In the optimization language, this corresponds to solving the inverse problem: for a given target value $f_{*}$, which NTFE FRSTs $\mathcal{T}$ give rise to decay constants $f(\mathcal{T},t_{\star}^{a})\approx f_{*}$? Below, we demonstrate that such problems can be addressed with our GA. For our search, we use $f_{*}=10^{14}$GeV as the target.
@@
$\text{NTFE FRSTs}\leq 4^{6}\times 64^{5}\times 168^{3}\times 734^{6}\approx 3.3\times 10^{36}\,.$ (4.10)
@@
Lastly, let us highlight the GA’s effectiveness in finding novel NTFE FRSTs compared to random sampling of two-face FRTs. In Fig. 8, we plot the number of “failed” extensions of configurations of two-face FRTs as a function of the generation. Let us reiterate that extensions “fail” whenever the FRTs of two-faces cannot be consistently extended to a regular triangulation of $\Delta^{\circ}$. At generation 0, we randomly sample FRTs of two-faces and need $4999$ choices $\mathcal{C}$ of two-face triangulations to obtain the desired $P=100$ NTFE FRSTs. As becomes evident from Fig. 8, the number of such unextandable choices of two-face triangulations decreases rapidly across the entire duration of the search. Said differently, the GA develops a strategy to find new configurations of two-face triangulations that give rise to novel FRSTs. In fact, over the entire space of states visited by the GA, only $8.4\%$ of FRSTs are redundant, i.e., appear more than once.
@@ 4.3 Maximizing axion-photon couplings — $h^{1,1}=491$
As a final application of our GA, we study the largest polytope in the KS database with $(h^{1,1},h^{1,2})=(491,11)$, whose associated Calabi-Yau geometries are expected to dominate the entire KS set. Indeed, according to [37], the search space of NTFE FRSTs of this polytope is bounded by [^26] The original bound of $\text{CYs}\lesssim 1.65\times 10^{428}$ obtained in [37] by estimating the number of two-face FRTs can be further improved to $\text{CYs}<10^{296.1}$ [82] by exactly counting and using more stringent bounds on the number of two-face FRTs of the polytope with $h^{1,1}=491$. $\text{CYs}\lesssim 1.65\times 10^{428}$. Because this polytope has both the largest search space in KS and features the most complex and computationally intensive geometries in this database, it represents an important and exciting target for optimization. Thanks to recent advances in computational geometry [6] rendering quantities such as triple intersection numbers tractable even at $h^{1,1}=491$, targeting this polytope with optimization techniques is now within reach.
@@
For the two-faces with $17,13,7$ and $3$ points, we find respectively $19594$, $204$, 55 and $1$ FRTs. There are three two-faces with more than $100$ points for which it is currently infeasible to compute the full set of FRTs. We explored two approaches to these two-faces: the simplest was merely to choose a Delaunay triangulation for each one, and allow the GA to explore the two-face triangulations for all other two-faces. We additionally investigated the use of random subsamples of the triangulations for these two-faces (using methods such as the ‘fast’ algorithm of [37]), but found that this decreased extendability substantially, so we decided to use the former approach and defer a more comprehensive study of the DNA at $h^{1,1}=491$ for future work. In summary, then, the GA’s search space had size
$\text{NTFE FRSTs accessible to GA}\leq 5^{2}\times 204^{2}\times 19594^{2}\approx 3.99\times 10^{14}\,.$ (4.13)
@@
Figure 9: Best $\|g_{a\gamma\gamma}\|$ found as a function of number of unique DNA encountered while optimizing, averaged over $17$ GA runs and compared with average performance of random DNA sampling. The average maximum value of $\log_{10}(\|g_{a\gamma\gamma}\|/\text{GeV}^{-1})$ after $5000$ unique evaluated DNA is $10.5\pm 0.22$ for the GA, and $11.3\pm 0.14$ for random sampling for $95\%$ confidence.
@@
Our optimization task is to maximize an effective overall axion photon coupling, given by the Euclidean norm of the vector of axion-photon couplings for each of the $h^{1,1}=491$ axions in our theory. We denote this $\|g_{a\gamma\gamma}\|$. For this to be well-defined, we must choose a divisor to host QED and a point in the Kähler cone: to pick these, we take some moderate inspiration from phenomenology. In particular, if we assume QCD and QED are hosted on separate divisors, the correct IR couplings are realized for divisor volumes of $40$ and $127.5$, respectively. We decide to compute the tip of the stretched Kahler cone and a basis of the smallest prime toric divisors, dilate until the smallest basis prime toric divisor $D$ has volume $40$ (thinking of this as our QCD divisor), and then place QED on the next smallest basis divisor $D^{\prime}$ which intersects $D$, requiring that this divisor have volume less than $127.5$. This determines our point in moduli space and the QED divisor: if such a dilation cannot be performed, or such a QED candidate divisor does not exist, we set the target to be -$\infty$. In this way, our $g_{a\gamma\gamma}$ computation is mildly phenomenologically informed, but we stress that our optimized values should taken with a grain of salt: we omit many crucial checks of the legitimacy of the associated low-energy EFT.
@@
We train GAs with target $\log_{10}\|g_{a\gamma\gamma}\|$ with population size $P=50$ until each has evaluated $5000$ unique DNA. It takes $\mathcal{O}(\text{hours})$ on standard hardware for a single run, rendering our hyperparameter optimization methods — for which we were performing hundreds of runs — impractical. We compare the outcome of the GA with random sampling with error bars from averaging over $10$ runs in Fig. 9. The GA distinguishes itself from a brute force search approach after exposure to only $\sim 500$ unique DNA from the search space and continues to improve at a much faster rate than random sampling. In particular, within $5000$ unique seen DNA the GA has found effective axion-photon couplings over an order of magnitude larger than encountered by random sampling. To make this more quantitative, the average maximum value of $\|g_{a\gamma\gamma}\|$ found by the two methods after $5000$ unique evaluated DNA is
@@
$\log_{10}(\|g_{a\gamma\gamma}\|/\text{GeV}^{-1})_{\text{max}}=\begin{cases}-10.5\pm 0.22&\text{Genetic Algorithm}\,,\\ -11.3\pm 0.14&\text{Random Sampling}\end{cases}$ (4.16)
@@
for $95\%$ confidence. Thus, we find the encouraging result that, even without any kind of tuning or hyperparameter optimization, a GA can improve upon brute-force searching even for the polytope with the largest and most computationally intricate search space.